Weekly LIPAc highlights (30/November-4/December)

Injector conditioning

The Injector operation continued this week: the duty cycle (dc) was increased up to 50% with a very stable beam and 100mA of total extracted current from deuterium plasma. The target performance for the facility inspection (50%dc @100mA) has been reached well ahead of schedule. Next week (last week of operation before the facility inspection, scheduled for 16th December) the efforts will concentrate on decreasing the ramp-up time (it currently takes a few hours to ramp-up the injector to 50% of duty cycle).

Screenshot of the injector oscilloscope on 03 December:

  • Duty cycle: 50% (80ms/160ms)
  • Total extracted current (red, Iext): ~100 mA
  • Yellow: Faraday cup signal (I_FC)
  • Blue: Current of repeller PS (I_repeller)
  • Green: Voltage of repeller PS (V_repeller)

Radio Frequency (RF) System and RFQ cavities (Radio Frequency Quadrupoles)

This week the RFQ nominal vacuum conditions were resumed switching on all the cryo pumping groups.

The entire RF system has been progressively started to validate the upgrades implemented in a part of the RF chains and to train QST personnel on the entire start-up process (calibration of the RF chains, RF-RFQ interfaces checks).  This afternoon we had successfully injected RF in the RFQ cavity!

Once completed the on-going tests and the upgrades validated, the modifications will be propagated in all the modules starting from next week, as soon as the RF system will be stopped. The target being to resume now the RF-RFQ conditioning in the second half of January.

HEBT (High Energy Beam Transport) commissioning

This week was very dense of remote sessions with great support from CIEMAT. The efforts focused on the commissioning and debugging of software for the different Beam Instrumentation equipment installed in the HEBT (SEM grid, Slits, Faraday Cup).

FMEA analyses

The Failure Mode and Effect Analysis (FMEA) of the LIPAc was initiated early this year with the objective to reduce hazards during its maintenance and operation phases. After 2 training sessions, the FMEA was discussed identifying hazards subsystem by subsystem during 26 working sessions. As a result, 59 actions and recommendations were proposed to improve the safety of the project and were included in the final report that was shared this week with the FMEA working group and group coordinators. The identified critical actions will be implemented as prerequisite before resuming the LIPAc beam operation. All the groups including the EU experts participated to this activity, so we can clearly call this achievement an all-hands work resulting from the collaboration of Japan and Europe within the Broader Approach.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (23-27/November)

Injector conditioning

The Injector operation continued this week: the duty cycle was increased up to 30% with a very stable beam and 100mA of total extracted current from deuterium plasma. The vault will now be predominantly closed to continue the conditioning of the Injector in preparation of the facility inspection. The target performances for the facility inspection (planned on December 16) is 50% duty cycle at 100 mA of total extracted current.

Screenshot of the injector oscilloscope on 25 November:

  • Duty cycle: 31.2% (50ms/160ms)
  • Red: Total extracted current (Iext)
  • Yellow: Faraday cup signal (I_FC)
  • Blue: Current of repeller PS (I_repeller)
  • Green: Voltage of repeller PS (V_repeller, -4.5 kV)

RadioFrequency (RF) System

The RF team concluded several important activities this week:

  • The issue on the High Voltage Power supply 3A (HVPS3A) was solved with continuous remote support from JEMA.
  • The upgrade of the PSYS (Protection system) interlock was completed and installed on one RF module. The modification of the threshold levels of PSYS VSWR (Voltage Standing Wave Ratio) card were completed on the same module.
  • The anode box upgrade to decrease the overheating at the circuit board was successfully implemented in one module.

The plan is to start next week the entire RF system to validate the upgrades implemented in a part of the RF chains and to train QST personnel on the entire start-up process. Once the upgrades validated, the RF system will be stopped and the upgrades extended to all the modules.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Remote data transfer success

A demonstration of data transfer from the Rokkasho BA site to F4E Garching was conducted on 23/10/2020 through a data diode… [See more].

Photo 1: Data transfer test at Central Control Room in Rokkasho
Photo 2: Data transfer test at F4E Garching

General LIPAc status of last week 45 (2-6/November)

Maintenance of Injector vacuum

The replacement of the oil rotary vane primary pump with a dry multi root pump was successfully completed… [See more]

Preparation for injector scroll pumps testing (using a new workbench installed in front of the injector to ease maintenance operation).

General current status and a nice technical achievement of last week 44 (26-30/October)

Maintenance of Injector vacuum

After the replacement of the turbomolecular pump last week, this week the replacement of the oil rotary vane primary pump with a dry multi root pump was performed… [See more]

Top view of the last section of the HEBT: from bottom to top, the doublet magnet, the dipole and the second triplet.

Project Committee #26

22 & 23 October 2020

The 26th Project Committee (PC#26) was held at Rokkasho IFERC Site on October 22nd and 23rd and by video conference… [See more]

General current status and a nice technical achievement of last week 43 (19-23/October)

Maintenance of Injector and Injector vacuum

The maintenance of the injector and its annexed LEBT (Low Energy Beam transport) was completed with the maintenance of the EMU … [See more]

General current status and a nice technical achievement of last week 42 (12-16/October)

Background information about general status

The LIPAC accelerator is currently shut down and in a planned long-term maintenance phase that is scheduled to be completed at the end of October… [See more]

The new RGBLM chamber after the completion of the installation in the MEL (the turbomolecular pump is placed horizontally and visible in the front)
Extended view of the MEL with the new vacuum chamber for the RGBLM

1st Fusion Neutron Source Technical Coordination Meeting

The first Fusion Neutron Source (FNS) technical coordination meeting was held on September 28-29 via videoconference… [See more]

IFERC and IFMIF/EVEDA Technical Coordination Meeting on remote participation

A joint IFERC/REC and IFMIF/EVEDA Technical Coordination Meeting was held on July 28-30 via videoconference… [See more]

Real-time information shared between QST and F4E

LIPAc maintenance progress

06 July 2020

LIPAc maintenance activities are on-going. On-sites activities have never stopped, but they were slowed down during a period of recommended teleworking due to Covid-19 outbreak… [See more]

Newly installed Medium-Energy Beam Transport extension line

More newsline….